OpenStudy (anonymous):

A dress manufacturer usually buys two rolls of cloth, one of 30 and 1/2 yards and the other of 29 and 3/4 yards, to fill his weekly orders. If his orders double one week, how much cloth should he order? (Give the total yardage). Enter your answer as a reduced mixed number or as a whole number

OpenStudy (anonymous):

30.5+29.75 = 60.25

OpenStudy (anonymous):

60.25 *2 = 120.5 120.5 = 120 and 1/2

OpenStudy (anonymous):

so which one is the answer?

OpenStudy (anonymous):

do you understand my first comment. (post what you think it is)

OpenStudy (anonymous):

is it 120 and 1/2

OpenStudy (anonymous):

first try to answer my question. it will hep you understand how I figured it out.

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

A dress manufacturer usually buys two rolls of cloth, one of 30 and 1/2 yards and the other of 29 and 3/4 yards the total amount he normally uses is 60.25 yards

OpenStudy (anonymous):

you need to double it. that's why i doubled it. 60.25*2 = 120.5 120.5 is 120 and 1/2 yards

OpenStudy (anonymous):

oh okay i see now